# app.build: A Production Framework for Scaling Agentic Prompt-to-App Generation with Environment Scaffolding

### v1

@@ 5.1 Environment Scaffolding Impact (TypeScript/tRPC only)
Table 3: Aggregated evaluation results for TypeScript/tRPC ($n=30$ prompts). Viability $V$ and quality $Q$ are defined in Section 4.5. “Perfect quality” denotes $Q=10$ (all applicable checks PASS). “Non-viable” denotes $V=0$ (AB-01 or AB-02 = FAIL). Mean quality is computed over viable apps only ($V=1$).
Metric Value Key Insight
Total Applications 30 TypeScript/tRPC stack only
Viability Rate ($V=1$) 73.3% 22/30 viable applications
Perfect Quality ($Q=10$) 30.0% 9/30 fully compliant applications
Non-viable ($V=0$) 26.7% 8/30 failed smoke tests
Mean Quality ($V=1$ apps) 8.78 High quality when viable
  • Note. Scoring rubric and check definitions in Section 4.5.
Table 4: Check-specific outcomes across $n=30$ TypeScript/tRPC tasks. See Section 4.5 for check definitions, PASS/WARN/FAIL grading, and the viability rule. NA indicates the check was not applicable to a prompt (e.g., AB-04 when no view/edit flows are required). “Pass Rate (excl. NA)” is computed over applicable cases only.
Check Pass Warn Fail NA Pass Rate (excl. NA)
AB-01 (Boot) 25 2 3 0 83.3%
AB-02 (Prompt) 19 3 5 3 70.4%
AB-03 (Create) 22 2 0 6 91.7%
AB-04 (View/Edit) 17 1 1 11 89.5%
AB-06 (Clickable Sweep) 20 4 1 5 80.0%
AB-07 (Performance) 23 3 0 4 88.5%
  • Note. AB-07 is a continuous metric normalized to $[0,1]$; thresholding for PASS/WARN/FAIL is specified in Section 4.5.
Smoke tests (AB-01, AB-02) determine viability. Among viable applications ($V=1$, $n=21$), quality averaged 8.78 with 77.3% achieving $Q\geq 9$. Non-viability ($V=0$) arises from smoke test failures or missing artifacts.
@@ 5.2 Open vs Closed Model Performance
We evaluated Claude Sonnet 4 against two open-weights models using the TypeScript/tRPC stack with simplified validation pipeline ensuring the app is bootable and renders correctly. Claude achieved 86.7% success rate, establishing our closed-model baseline at \$110.20 total cost. Qwen3-Coder-480B-A35B reached 70% success rate (80.8% relative performance) while GPT OSS 120B managed only 30% success rate. Both open models were accessed via OpenRouter, resulting in significantly lower costs: \$12.68 for Qwen3 and \$4.55 for GPT OSS.
The performance gap reveals that environment scaffolding alone cannot eliminate the need for capable foundation models. However, leading open-weights models like Qwen3 demonstrate that structured environments can enable production-viable performance at substantially reduced costs. The 9x cost reduction for 19% performance loss represents a viable tradeoff.
Operational characteristics differed notably between model types. Open models required more validation retries, evidenced by higher LLM call counts (4,359 for Qwen3, 4,922 for GPT OSS vs 3,413 for Claude). Healthcheck pass rates (86.7% for Qwen3 vs 96.7% for Claude) indicate open models generate syntactically correct code but struggle with integration-level correctness, emphasizing the importance of comprehensive validation.
@@ 5.3 Ablation Studies: Impact of Validation Layers
To understand how each validation layer contributes to application quality, we conducted controlled ablations on the same 30-prompt cohort. Each ablation removes one validation component while keeping others intact.
Baseline Performance (all validation layers active):
  • Viability: 73.3% (22/30 apps pass both AB-01 Boot and AB-02 Prompt)
  • Mean Quality: 8.06 (among all 30 apps)
Finding 1: Removing Unit Tests Trades Quality for Viability
  • Viability: 80.0% (+6.7 pp) – fewer apps fail smoke tests
  • Mean Quality: 7.78 (0.28-0.28) – quality degrades despite higher viability
  • Key degradations: AB-04 View/Edit drops from 90% to 60% pass rate
  • Interpretation: Backend tests catch critical CRUD errors. Without them, apps boot successfully but fail on data operations.
Finding 2: Removing Linting Has Mixed Effects
  • Viability: 80.0% (+6.7 pp)
  • Mean Quality: 8.25 (+0.19) – slight improvement
  • Trade-offs: AB-03 Create drops 8.3 pp, AB-04 View/Edit drops 7.6 pp
  • Interpretation: ESLint catches legitimate issues but may also block valid patterns. The performance gain suggests some lint rules may be overly restrictive.
Finding 3: Removing Playwright Tests Significantly Improves Outcomes
  • Viability: 90.0% (+16.7 pp) – highest among all configurations
  • Mean Quality: 8.62 (+0.56) – meaningful quality improvement
  • Broad improvements: AB-02 Prompt +11.8 pp, AB-06 Clickable +5.7 pp
  • Interpretation: Playwright tests appear overly brittle for scaffolded apps. Many apps that fail E2E tests actually work correctly for users.
@@ 5.4 Synthesis: Optimal Validation Strategy
Our ablation results reveal clear trade-offs in validation design:
Validation Layer Impact Summary:
  1. 1.
    Unit/Handler Tests: Essential for data integrity. Removing them increases perceived viability but causes real functional regressions (especially AB-04 View/Edit).
  2. 2.
    ESLint: Provides modest value with some false positives. The small quality impact (+0.19) and mixed per-dimension effects suggest selective application.
  3. 3.
    Playwright/E2E: Currently causes more harm than good. The +16.7 pp viability gain and quality improvements indicate these tests reject too many working applications.
Recommended Validation Architecture: Based on these findings, we recommend:
  • Keep: Lightweight smoke tests (boot + primary route), backend unit tests for CRUD operations
  • Refine: ESLint with curated rules focusing on actual errors vs style preferences
  • Replace: Full E2E suite with targeted integration tests for critical paths only
This pragmatic approach balances catching real defects while avoiding false rejections. When quality is paramount and compute budget less constrained, comprehensive validation including strict E2E tests remains viable—trading lower success rates for guaranteed production quality.
@@ 5.5 Failure Mode Analysis
Failure modes in tRPC runs cluster into categories:
  • Boot/Load failures: template placeholders or incomplete artifacts
  • Prompt correspondence failures: generic templates from generation failures
  • CSP/security policy restrictions: blocked images or media by default policies
  • UI interaction defects: unbound handlers, non-working controls
  • State/integration defects: data not persisting across refresh; broken filters; login issues
  • Component misuse: runtime exceptions from incorrect component composition
These defects align with our layered pipeline design: early gates catch non-viable builds, while later gates expose interaction/state issues before human evaluation.
@@ 5.6 Prompt Complexity and Success Rate
We categorize prompts along a simple rubric and analyze success impacts:
  • Low complexity: static or single-page UI tasks (e.g., landing pages, counters)
  • Medium complexity: single-entity CRUD without advanced flows or auth
  • High complexity: multi-entity workflows, custom logic, or complex UI interactions
@@
Medium-complexity CRUD prompts achieve the highest quality (Q=9Q=9–10), reflecting strong scaffolding for data models and handlers. Low-complexity UI prompts are not uniformly easy: several failed prompt correspondence (AB-02) with generic templates. High-complexity prompts show lower viability rates due to interaction wiring and state-consistency issues surfaced by AB-04/AB-06.
6 Discussion
6.1 Limitations
@@
Our current framework is limited to CRUD-oriented data applications, focusing on structured workflows with well-defined input-output expectations. While effective for common web application patterns, it does not yet support complex systems or advanced integrations. The validation pipeline, though comprehensive, relies on domain-specific heuristics and expert-defined anti-patterns, which may not generalize to novel or edge-case designs. Additionally, our human evaluation protocol, while rigorous, is poorly scalable and constrained by subjectivity in assessing maintainability and user experience nuances.
6.2 Broader Impact
@@
7 Conclusion
Our results demonstrate that raw model capability alone cannot bridge the gap between AI potential and production reality. Through systematic environment scaffolding, multi-layered validation, and stack-specific orchestration, app.build transforms probabilistic language models into dependable software engineering agents.
@@ 7 Conclusion
The path to reliable AI agents lies not in better prompts or bigger models, but in principled environment engineering with validation layers tuned to maximize value while minimizing brittleness.
8 Acknowledgments
@@
Appendix A Prompt Dataset (Full List)
Table 5: Complete prompt dataset used in evaluation ($n=30$). Dataset construction details in Section 4.2. Complexity labels follow the rubric in Section 5.6: Low (static/single-page UI), Medium (single-entity CRUD), High (multi-entity/custom logic).
@@ Appendix A Prompt Dataset (Full List)
kanji-flashcards Kanji learning with SRS, progress tracking, JLPT levels. High

### v2

@@ 5.5 Detailed Quality Assessment (Human Evaluation, n=30)
Table 5: Aggregated Evaluation Results for TypeScript/tRPC ($n=30$)
Metric Value Note
Total Apps 30 tRPC stack only
Viability ($V=1$) 73.3% 22/30 viable
Perfect ($Q=10$) 30.0% 9/30 perfect
Non-viable ($V=0$) 26.7% 8/30 failed
Mean Quality 8.78 $V=1$ apps only
  • Viability $V$ and quality $Q$ defined in Section 4.4. Perfect = all checks PASS; non-viable = AB-01 (Boot) or AB-02 (Prompt Correspondence) FAIL.
Table 6: Check-Specific Outcomes Across $n=30$ Tasks
Check Pass Warn Fail NA
AB-01 (Boot) 25 2 3 0
AB-02 (Prompt Correspondence) 19 3 5 3
AB-03 (Create Functionality) 22 2 0 6
AB-04 (View/Edit Operations) 17 1 1 11
AB-05 (Clickable Sweep) 20 4 1 5
AB-06 (Performance) 23 3 0 4
  • See Section 4.4 for grading criteria. NA = not applicable for prompt type. Pass rates (excl. NA): AB-01 (Boot): 83.3%, AB-02 (Prompt Correspondence): 70.4%, AB-03 (Create Functionality): 91.7%, AB-04 (View/Edit Operations): 89.5%, AB-05 (Clickable Sweep): 80.0%, AB-06 (Performance): 88.5%.
Smoke tests (AB-01, AB-02) determine viability. Among viable applications ($V=1$, $n=22$), quality averaged 8.78 with 77.3% achieving $Q\geq 9$. Non-viability ($V=0$) arises from smoke test failures or missing artifacts.
@@ 5.6 Open vs Closed Model Performance
We evaluated Claude Sonnet 4 against two open-weights models using the TypeScript/tRPC stack with simplified validation pipeline (AB-01 Boot + AB-02 Prompt Correspondence only) ensuring the app is bootable and renders correctly. Claude achieved 86.7% success rate, establishing our closed-model baseline at \$110.20 total cost. Qwen3-Coder-480B-A35B reached 70% success rate (80.8% relative performance) while GPT OSS 120B managed only 30% success rate. Both open models were accessed via OpenRouter, resulting in significantly lower costs: \$12.68 for Qwen3 and \$4.55 for GPT OSS.
The performance gap reveals that environment scaffolding alone cannot eliminate the need for capable foundation models. However, leading open-weights models like Qwen3 demonstrate that structured environments can enable production-viable performance at substantially reduced costs. The 8.2x cost reduction per viable app for 19% performance loss represents a viable tradeoff for many production scenarios.
Template detection reveals inflated success rates. Our evaluation uncovered a critical issue: many apps that pass healthcheck validation are non-functional "Under Construction" templates with zero functionality. For GPT-OSS-120B, 43.3% passed AB-01 (Boot), but only 30.0% passed both AB-01 and AB-02 (Prompt Correspondence), with manual inspection revealing that a substantial portion of bootable apps were generic placeholder scaffolding rather than functional implementations. In contrast, Qwen3-Coder-480B-A35B achieved 86.7% healthcheck pass and 70.0% true success, showing better prompt adherence with fewer template generations. The Success column in Table 3 reports only viable non-template apps that passed AB-02 validation, while HC Pass includes all bootable apps regardless of functionality, demonstrating why boot checks alone are insufficient for evaluating code generation quality.
Operational characteristics differed notably between model types. Open models required more validation retries, evidenced by higher LLM call counts (4,359 for Qwen3, 4,922 for GPT OSS vs 3,413 for Claude). AB-01 (Boot) pass rates (86.7% for Qwen3 vs 96.7% for Claude) indicate open models generate syntactically correct code but struggle with integration-level correctness, emphasizing the importance of comprehensive validation.
@@ 5.7 Ablation Studies: Impact of Validation Layers
To understand how each validation layer contributes to application quality, we conducted controlled ablations on the same 30-prompt cohort, systematically removing one validation component while keeping others intact. The baseline configuration (all validation layers active) achieved 73.3% viability with mean quality score $Q=8.06$.
Unit test removal trades quality for apparent viability. Disabling backend handler tests increased viability to 80.0% (+6.7 pp) but reduced mean quality to $Q=7.78$ (0.28-0.28). This paradox reflects that unit tests catch critical CRUD errors: apps boot successfully without them but fail on data operations. AB-04 (View/Edit Operations) pass rates dropped from 90% to 60%, indicating that backend validation prevents functional regressions that smoke tests cannot detect. The increased viability metric reflects false positives: apps that appear viable but contain latent data integrity bugs.
Linting removal shows mixed effects with modest gains. Removing ESLint checks increased viability to 80.0% (+6.7 pp) and slightly improved quality to $Q=8.25$ (+0.19), suggesting some lint rules may be overly restrictive. However, AB-03 (Create Functionality) dropped 8.3 pp and AB-04 (View/Edit) dropped 7.6 pp, indicating ESLint does catch legitimate structural issues. The net positive quality score indicates that strict linting can reject valid alternative implementations, though the effect size is small.
E2E test removal significantly improves outcomes (a negative result). Removing Playwright tests produced the strongest effect: viability increased to 90.0% (+16.7 pp) with quality improving to $Q=8.62$ (+0.56). AB-02 (Prompt Correspondence) improved +11.8 pp and AB-05 (Clickable Sweep) improved +5.7 pp. This counterintuitive result indicates that E2E tests introduce more false rejections than they catch real defects. Manual inspection of failed Playwright runs reveals three root causes of brittleness: (1) Over-specified selectors: tests hardcode element IDs or CSS classes that models generate variably across runs, causing spurious failures despite functional correctness; (2) Race conditions: E2E assertions check UI state before async data fetching completes, rejecting apps that function correctly under human-realistic interaction delays; (3) False negatives from implementation variance: models generate semantically equivalent but structurally different UIs (e.g., modal dialogs vs inline forms) that satisfy requirements but fail brittle assertions expecting specific Document Object Model (DOM) structure. This finding suggests that comprehensive E2E suites designed for deterministic codebases are poorly suited to probabilistic code generation, where implementation details vary across generation runs while functional correctness remains stable.
@@ 5.8 Synthesis: Optimal Validation Strategy
Our ablation results reveal systematic trade-offs between validation rigor and success metrics. Unit/handler tests prove essential for data integrity: removing them increases perceived viability (+6.7 pp) but causes real functional regressions, particularly in AB-04 (View/Edit Operations, 30-30 pp drop). ESLint provides modest value with measurable false positives; the small net quality gain (+0.19) and mixed per-check effects suggest selective application of rules targeting actual errors rather than style preferences. E2E tests currently cause more harm than good, with removal yielding +16.7 pp viability and +0.56 quality improvement, indicating these tests reject too many working applications due to brittleness rather than correctness issues.
These findings suggest a pragmatic validation architecture optimized for probabilistic code generation: retain lightweight smoke tests (boot verification and primary route checks) combined with backend unit tests for CRUD operations, which together provide high-confidence quality gates; refine static analysis to focus on structural correctness and known anti-patterns while relaxing stylistic rules that reject valid implementations; replace comprehensive E2E suites with targeted integration tests covering only critical user paths, avoiding brittle assertions on implementation details. This configuration balances defect detection against false rejection rates. For scenarios where quality requirements dominate cost constraints, comprehensive validation including strict E2E tests remains viable, trading lower automated success rates (16.7-16.7 pp) for stronger guarantees about production correctness.
@@ 5.9 Failure Mode Analysis
We systematically categorize observed failure modes and map them to AB check coverage. Boot/Load failures (template placeholders, incomplete artifacts) are reliably caught by AB-01 (Boot) during server startup. Prompt correspondence failures (generic templates from generation failures) are detected by AB-02 when human evaluators identify that apps render successfully but implement incorrect functionality. Content Security Policy (CSP) restrictions (blocked images or media) are partially caught by AB-05 (Clickable Sweep) through visual inspection, though automated checks miss these unless they break critical flows. UI interaction defects (unbound event handlers, non-functional controls) are exposed by AB-05 and AB-03 (Create Functionality) through manual interaction; automated E2E tests theoretically detect these but suffer from brittleness as documented in Section G. State/integration defects (data persistence failures across refresh, broken filters, authentication issues) are caught by AB-04 (View/Edit Operations), though backend unit tests provide only partial coverage and miss client-side state management bugs. Finally, component misuse (runtime exceptions from incorrect composition) is detected by AB-01 or AB-05 depending on timing, with TypeScript catching some violations at compile time while others surface only during manual testing.
Validation blind spots and proposed upgrades. Our AB check suite exhibits systematic coverage gaps across five dimensions. First, accessibility violations go undetected; apps may function for able-bodied users but fail Web Content Accessibility Guidelines (WCAG) standards. We propose integrating axe-core for automated accessibility auditing. Second, mobile responsiveness remains unchecked as all validation uses desktop viewports; extending smoke tests to include mobile viewport validation would address this. Third, performance degradation under load is not measured; AB-06 assesses only initial load time, missing runtime performance issues. Lighthouse performance budgets could provide continuous monitoring. Fourth, security vulnerabilities beyond CSP violations (Cross-Site Scripting (XSS), Cross-Site Request Forgery (CSRF), injection attacks) are not detected; static analysis via semgrep with OWASP rulesets would provide defense-in-depth. Fifth, data consistency in concurrent scenarios and transaction isolation remain untested; implementing multi-client test scenarios would validate state consistency guarantees. These proposed upgrades would strengthen production readiness while maintaining the environment scaffolding paradigm of deterministic, automated validation.
@@ 5.10 Prompt Complexity and Success Rate
We categorize prompts by complexity: low (static/single-page UI), medium (single-entity CRUD), high (multi-entity workflows with custom logic), and analyze success patterns. Medium-complexity CRUD prompts achieve highest quality (Q=9Q=9–10), reflecting strong scaffolding alignment with data models and handler patterns. Low-complexity UI prompts prove non-uniformly easy: several failed AB-02 (Prompt Correspondence) by generating generic templates rather than task-specific implementations. High-complexity prompts exhibit lower viability due to interaction wiring and state-consistency issues surfaced by AB-04 (View/Edit Operations) and AB-05 (Clickable Sweep).
Trajectory patterns reveal scaffolding trade-offs. Manual inspection of generation logs exposes distinct model reasoning patterns across complexity levels. For medium-complexity CRUD tasks, environment scaffolding provides effective guidance: structured schema definitions constrain data models to valid patterns, type-safe API contracts prevent interface mismatches, and immediate validator feedback enables targeted repairs within 1–2 iterations. Representative trajectory (plant-care-tracker): schema generation passed immediately; API handlers failed unit tests due to missing foreign key constraints, triggering repair with validator feedback; UI generation succeeded on first attempt using type inference from validated API. This demonstrates the intended scaffolding workflow.
Conversely, for low-complexity UI-only tasks, scaffolding introduces bias: models over-engineer solutions using CRUD templates when simple static pages suffice, causing AB-02 failures. The framework’s database-first workflow assumes backend requirements even when unnecessary. Failure example (birthday-wish-app): model generated full API routes and database schema for a static greeting card, achieving technical correctness but failing prompt correspondence. This reveals a structural limitation where strong scaffolding can constrain creativity for tasks outside the target domain.
High-complexity tasks expose scaffolding’s context limits. Multi-entity relationships and custom business logic require models to maintain architectural context across multiple validation cycles spanning frontend-backend boundaries. Failures cluster in AB-04 (state management) when models lose track of data flow requirements. Validation feedback successfully catches syntax errors but provides limited architectural guidance, forcing models into trial-and-error that exhausts retry budgets before convergence. This suggests that scaffolding effectiveness degrades as task complexity increases beyond the structured patterns the environment encodes.
5.11 Threats to Validity & Limitations
@@
Our current framework is limited to CRUD-oriented data applications, focusing on structured workflows with well-defined input-output expectations. While effective for common web application patterns, it does not yet support complex systems or advanced integrations. The evaluation focused exclusively on the TypeScript/tRPC stack because it is the most mature and well-supported stack in our framework, enabling the most reliable assessment of environment scaffolding effectiveness. We chose web application generation as the evaluation domain because it represents the most widely adopted use case among both closed-source commercial generators and academic end-to-end code generation systems, providing a relevant benchmark for comparison. The validation pipeline, though comprehensive, relies on domain-specific heuristics and expert-defined anti-patterns, which may not generalize to novel or edge-case designs. Additionally, our human evaluation protocol, while rigorous, is poorly scalable and constrained by subjectivity in assessing maintainability and user experience nuances.
Benchmark applicability. We do not evaluate on SWE-bench (Jimenez et al., 2024) or HumanEval (Chen et al., 2021) because these benchmarks target fundamentally different tasks: repository-level bug fixing and function-level code completion, respectively. SWE-bench evaluates patch correctness against existing test suites in mature codebases, while our work generates complete applications requiring multi-layered validation (schema correctness, API contracts, UI functionality, integration testing). The evaluation methodology mismatch is structural: existing benchmarks assume deterministic test oracles, whereas greenfield application generation requires human assessment of prompt correspondence, as our template detection findings demonstrate—53.6% of GPT-OSS-120B outputs passed automated healthchecks but were non-functional placeholders. Our ablation studies provide baseline comparisons that demonstrate environment scaffolding’s value: configurations with validation layers disabled show that removing backend tests increases apparent viability (+6.7pp) but causes CRUD correctness to drop 30pp, revealing that comprehensive validation trades automated success metrics for functional correctness.
5.12 Ethics & Broader Impact
@@
Our results suggest that for CRUD-oriented web applications, structured environment scaffolding complements model capability in achieving production reliability. Through systematic validation, stack-specific orchestration, and iterative repair, app.build demonstrates how probabilistic language models can be guided toward dependable software generation within constrained domains.
@@
For production-oriented agent systems in structured domains, environment engineering with targeted validation layers offers a complementary path to scaling model capability, providing measurable improvements in reliability while managing cost. As model capabilities continue to advance, the systematic integration of validation and iterative repair remains essential for bridging the gap between probabilistic generation and deterministic production requirements.
Data and Artifact Availability
To support reproducibility, we release the following artifacts:
Acknowledgments
@@
Appendix A Prompt Dataset
This appendix provides the complete specification of evaluation prompts used in our experiments. Table 7 lists all 30 prompts with their complexity classifications, covering low-complexity UI tasks, medium-complexity single-entity CRUD applications, and high-complexity multi-entity workflows with custom business logic.
Table 7: Complete Prompt Dataset Used in Evaluation ($n=30$)
@@ Appendix A Prompt Dataset
kanji-flashcards Kanji learning with Spaced Repetition System (SRS), progress tracking, JLPT levels. High
@@
Note. Dataset details in Section 4.2. Complexity rubric in Section 5.10: Low (static/single-page UI), Medium (single-entity CRUD), High (multi-entity/custom logic).